The yield to maturity on 1-year zero-coupon bonds is currently 4.5%; the YTM on 2-year zeros is 5.5%. The Treasury plans to issue a 2-year maturity coupon bond, paying coupons once per year with a coupon rate of 6% The face value of the bond is $100. a. At what price will the bond sell? (Do not round intermediate calculations. Round your answer to 2 decimal places.) Price b. What will the yield to maturity on the bond be? (Do not round intermediate calculations. Round your answer to 3 decimal places.) Yield to maturity c. If the expectations theory of the yield curve is correct, what is the market expectation of the price that the bond will sell for next year? (Do not round intermediate calculations. Round your answer to 2 decimal places.) Price d. Recalculate your answer to (c) if you believe in the liquidity preference theory and you believe that the liquidity premium is 1%. (Do not round intermediate calculations. Round your answer to 2 decimal places.) Price
Answer #1
year CF PV factor PV
1 6 0.956938 5.7416268
2 106 0.898452 95.235956
Price $     100.98

b)=RATE(2,6,-100.98,100)
=5.47%

c)1 year rate next year = 1.055^2/1.045= 6.5096%



Value of bond = 106/1.065096=$99.52

d) Addding up 1% rate = 7.5096%
Market expectation pric eof bond= 106/1.075096=$98.60